OpenStudy (anonymous):

find d^2y/dx^2 of y^2=x^2+2x

OpenStudy (anonymous):

find y' first

OpenStudy (anonymous):

\[2yy'=2x+2\] \[y'=\frac{2x+2}{2y}=\frac{x+1}{y}\]

OpenStudy (anonymous):

then \[y''=\frac{y-(x+1)y'}{y^2}\] and now replace y' by what you know it to be

OpenStudy (anonymous):

how did you get \[y \prime \prime \]?

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

okay i got that, then what?

OpenStudy (anonymous):

\[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x+1,f'(x)=1,g(x)=y, g'(x)=y'\]

OpenStudy (anonymous):

now some annoying algebra. you get \[\frac{y-(x+1)\times \frac{x+1}{y}}{y^2}\] multiply top and bottom by y to clear the compound fraction get \[\frac{y^2-(x+1)^2}{y^3}\]

OpenStudy (anonymous):

and then you are still not done.

OpenStudy (anonymous):

because if you multiply out in the numerator you get \[y^2-(x^2+2x+1)=y^2-y^2-1=-1\]

OpenStudy (anonymous):

so your "final answer" i think is \[y''=-\frac{1}{y^3}\]

OpenStudy (anonymous):

how did you get the second \[y ^{2}\]

OpenStudy (anonymous):

in \[y ^{2}-y ^{2} -1\]

OpenStudy (anonymous):

your original equation was \[y^2=x^2+2x\] so replaced \[x^2+2x\] by \[y^2\]

OpenStudy (anonymous):

OMG!!! thats some hard shxt

OpenStudy (anonymous):

oh okay that makes sense thank you so much!

OpenStudy (anonymous):

there is another way to do it but it is not really any easier, always look to the original equation when you are done, because often it is cooked up to simplify a bunch

OpenStudy (anonymous):

yw